A 1967 Corvette of mass 1450 kg moving with a velocity of 100 mph (= 44.7 m/s) slides on a slick street and collides with a Hummer of mass 3250 kg which is parked on the side of the street. The two vehicles interlock and slide off together. What is the speed of the two vehicles immediately after they join?
1
Expert's answer
2019-10-23T09:42:50-0400

The law of conservation of momentum gives


"mv=(m+M)u"

So


"u=\frac{mv}{m+M}=\frac{1450\times 44.7}{1450+3250}=13.8\:\rm m/s=29\: mph"